# Unifying Simulation and Inference with Normalizing Flows

### v2

@@ I Introduction
Machine learning provides a set of tools that can analyze hadronic final states holistically to achieve the best precision. Recent examples supporting this claim include the energy calibrations of single hadrons [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], jets [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], and global event properties [22, 23, 24, 25] at colliders. These techniques can automatically make use of finer segmentation to improve the resolution of reconstructed particle energies [4, 10]. While machine learning has broader utility than reconstructing individual hadron showers within a calorimeter, we focus on this case because it is particularly challenging and a critical component of event reconstruction at the Large Hadron Collider and elsewhere.
@@
Over the last years, some solutions to these two central challenges have been proposed. Generalized numerical inversion [26, 27, 28] is a prior-independent calibration approach, although it does not scale well to many output dimensions. The Gaussian Ansatz [17, 18] is a maximum likelihood estimator, and thus prior-independent (more in Sec. II). For UQ, loss functions can be modified to estimate quantiles in addition to just the mean, median or mode [13, 14] and deep generative models can estimate the exact resolution function [24, 25]. The Gaussian Ansatz produces an estimate of the resolution and so is currently the only method that is both prior-independent and that goes beyond a point estimate. We note that it is also possible to use Bayesian versions [29, 30, 31, 32, 33] of neural networks to go beyond point estimates by providing the uncertainty of the output. However, the uncertainty would still be prior-dependent.
@@ II.2 Maximum Likelihood Inference
Although the MLE-based calibration is always prior-independent, the prior dependence does not guarantee that the method is always unbiased. Only in certain scenarios (e.g. 1D Gaussian noise model such that $p_{X|Z}(x|z)\sim N(z,\sigma)$) can the MLE calibration be shown to be unbiased. Certain detector responses $p_{X|Z}$ may result in biased calibration. In general, prior independence is a necessary but insufficient condition for unbiased calibration. Even so, we show in Sec. III.3 that MLE-based calibration results in a smaller bias compared to direct regression for an ATLAS-like calorimeter setup. This suggests that in our application, the dependence on training prior is a bigger contributing factor to the bias than the detector response.
@@ III.3 Prior-independent and less biased calibration
    Draw with replacement $N$ samples from $N$ values of $E_{\rm pred}$, where $N$ is the number of showers in the evaluation dataset for a given fixed $E_{\rm true}$.
@@
    Repeat steps 1-3 for a total of 20 times.
@@
    Compute the mean and standard deviation of the 20 estimated values of the mode.
Figure 2: Plot of likelihood evaluated using CaloFlow for a given calorimeter shower originating from a 30 GeV pion (left), 60 GeV pion (middle) and 90 GeV pion (right), respectively. In each plot, the true incident energy $E_{\rm true}$ is shown by the vertical black line, and the evaluated likelihood by the NF is shown in blue. The lower (upper) bounds of the 68% confidence interval about the MLE is shown by the red vertical dashed lines.
The mean of the multiple mode estimates is taken to be the final mode estimate at each fixed $E_{\rm true}$. The corresponding standard deviation is taken to be the uncertainty of the final mode estimate. The mode estimation is performed for both the uniform and log-uniform cases. In Fig. 3, we show the the mode of $E_{\rm pred}/E_{\rm true}$ distribution at $E_{\rm true}\in\{10,20,30,40,50,60,70,80,90\}$ GeV for the different regression methods. Here we used nine evaluation datasets, each with 100k $\pi^{+}$ showers at one of the nine fixed $E_{\rm true}$ values.
@@
We observe that the flow-based calibration is generally less biased (closer to unity) compared to the MSE-based calibration. The larger bias of the MSE-based calibration is clearly seen at the edges of $E_{\rm true}$ range. Furthermore, we see that the MSE-based calibration corresponding to the log-uniform prior is generally more biased than the calibration corresponding the uniform prior at intermediate energies.
@@ III.4 Resolution estimation
Although not shown in Fig. 4, we found that the MSE-based calibration obtains a full resolution that closely matches that from the NF. However, it is unable to provide per-shower resolutions.
@@
Since the NF is able to access the complete likelihood, this enables it to capture non-Gaussian [^8] In principle, the Gaussian Ansatz [17, 18] should also be able to capture non-Gaussian components of the resolution. However, this has yet to be shown in a worked example. Previous works [18, 91] only attempted to compute the Gaussian resolution. Also, Gaussian Ansatz is trained in a different way (see Ref. [17]) compared to the NF and is more sensitive to hyperparameter choices. components of the resolution. We quantify this by defining an asymmetry observable based on the upper and lower bounds of the 68% confidence interval of the MLE prediction. In particular, the asymmetry is computed as the ratio $\frac{\Delta_{L}}{\Delta_{U}}$, where $\Delta_{L}$ ($\Delta_{U}$) is the difference between the lower (upper) bound of the 68% confidence interval and the MLE.
@@ IV Conclusions
We demonstrate a second advantage of using a MLE-based calibration methods over direct regression by estimating per-shower resolutions. The average estimated per-shower resolution obtained from CaloFlow closely aligns with the full resolution. In contrast, the direct regression approach yields only point estimates and lacks the capacity to make per-shower resolution predictions. For CaloFlow, we found that smaller per-shower resolutions tend to coincide with more accurately calibrated showers. This is evidence that the per-shower resolutions from CaloFlow are reliable and have the potential to be leveraged for improved calibration. Notably, utilizing a NF for calibration grants access to the complete resolution function. This characteristic enables the NF to capture asymmetries in the resolution. In future work, it would be interesting to study how the per-shower resolution information obtained by the NF can be utilized to further improve the calibration.
@@ Appendix A Architecture and training
Here we briefly describe the architecture and training procedure used for CaloFlow (see Refs. [42, 43] for more details). There are some differences compared to the implementation in the original CaloFlow papers [42, 43], but most of main algorithm remains the same. One main difference is that only Flow-I is used in this study.
The flows used in this work are Masked Autoregressive Flows (MAFs) [93] with compositions of Rational Quadratic Splines (RQS) [94] as transformations. The RQS transformations are parameterized using neural networks known as MADE blocks [95]. Identical flow architectures are used in each of the two cases with uniformly and log-uniformly distributed $E_{\rm inc}$. Each flow consists of six MADE blocks, each with two hidden layers of 64 nodes. The RQS transformations are defined with 8 bins and a tail bound of 14.

### v3

@@ I Introduction
Machine learning provides a set of tools that can analyze hadronic final states holistically to achieve the best precision. Recent examples supporting this claim include the energy calibrations of single hadrons [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], jets [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], and global event properties [22, 23, 24, 25] at colliders. These techniques can automatically make use of finer segmentation to improve the resolution of reconstructed particle energies [4, 10]. While machine learning has broader utility than reconstructing individual hadron showers within a calorimeter, we focus on this case because it is particularly challenging and a critical component of event reconstruction at the Large Hadron Collider and elsewhere. The process of inverting detector simulation is closely related to unfolding. Various unfolding methods have been developed and are actively utilized in experimental research. For comprehensive reviews, refer to Refs. [26, 27, 28, 29], and for the most commonly used unfolding algorithms, see Refs. [30, 31, 32]. The main difference between unfolding and calibration is that unfolding aims to correct distributions, while calibration aims to correct individual objects/events. Recently, there have been several proposed ML-based unfolding methods that are able to achieve remarkable unfolding performance. See Refs. [33, 34, 35, 36, 37, 38, 39, 24, 40, 41, 42, 43, 44, 45, 46] for examples of ML-based unfolding methods.
@@
Over the last years, some solutions to these two central challenges have been proposed. Generalized numerical inversion [47, 48, 49] is a prior-independent calibration approach, although it does not scale well to many output dimensions. The Gaussian Ansatz [17, 18] is a maximum likelihood estimator, and thus prior-independent (more in Sec. II). For UQ, loss functions can be modified to estimate quantiles in addition to just the mean, median or mode [13, 14] and deep generative models can estimate the exact resolution function [24, 25]. In the CMS Run 1 Higgs analysis [50], the uncertainty of the photon energy regressed by a boosted decision tree (BDT) is estimated by training an additional BDT. The Gaussian Ansatz produces an estimate of the resolution and so is currently the only method that is both prior-independent and that goes beyond a point estimate. We note that it is also possible to use Bayesian versions [51, 52, 53, 54, 55] of neural networks to go beyond point estimates by providing the uncertainty of the output. However, the uncertainty would still be prior-dependent.
@@ II.2 Maximum Likelihood Inference
Although the MLE-based calibration is always prior-independent, the prior independence does not guarantee that the method is always unbiased. Only in certain scenarios (e.g. 1D Gaussian noise model such that $p_{X|Z}(x|z)\sim N(z,\sigma)$) can the MLE calibration be shown to be unbiased. Certain detector responses $p_{X|Z}$ may result in biased calibration. In general, prior independence is a necessary but insufficient condition for unbiased calibration. Even so, we show in Sec. III.3 that MLE-based calibration results in a smaller bias compared to direct regression for an ATLAS-like calorimeter setup. This suggests that in our application, the dependence on training prior is a bigger contributing factor to the bias than the detector response.
@@ III.3 Prior-independent and less biased calibration
    Obtain predicted energies:
    For a given $E_{\rm true}$ (e.g., 30 GeV), collect $N$ values of $E_{\rm pred}$ predicted by the model. Here $N$ is the number of showers in the evaluation dataset for a given fixed $E_{\rm true}$. For each $E_{\rm pred}$ value, form the ratio $E_{\rm pred}/E_{\rm true}$.
@@
    Bootstrap resampling:
    Draw with replacement $N$ samples from $N$ values of $E_{\rm pred}$.
  3. 3.
    Kernel density estimation (KDE):
@@
    Identify the mode:
    Identify the position of the mode of the estimated density. This value represents the mode of the distribution for that particular bootstrap sample.
@@
    Repeat and accumulate:
    Repeat steps 2-4 for a total of 20 times. Store the mode estimate from each iteration.
  6. 6.
    Compute final estimate and uncertainty:
    • Final mode estimate: Take the mean of the 20 mode values obtained in the previous step.
    • Uncertainty: Use the standard deviation of those 20 mode estimates as the uncertainty on the final mode.
Figure 2: Plot of likelihood evaluated using CaloFlow for a given calorimeter shower originating from a 30 GeV pion (left), 60 GeV pion (middle) and 90 GeV pion (right), respectively. In each plot, the true incident energy $E_{\rm true}$ is shown by the vertical black line, and the evaluated likelihood by the NF is shown in blue. The lower (upper) bounds of the 68% confidence interval about the MLE are shown by the red vertical dashed lines.
The mode estimation is performed for both the uniform and log-uniform cases. In Fig. 3, we show the the mode of $E_{\rm pred}/E_{\rm true}$ distribution at $E_{\rm true}\in\{10,20,30,40,50,60,70,80,90\}$ GeV for the different regression methods. Here we used nine evaluation datasets, each with 100k $\pi^{+}$ showers at one of the nine fixed $E_{\rm true}$ values.
@@
We observe that the flow-based calibration is generally less biased (closer to unity) compared to the MSE-based calibration. The larger bias of the MSE-based calibration is clearly seen at the edges of the $E_{\rm true}$ range. We note that it may be possible to address this limitation with e.g. domain continuation [110]. Domain continuation has been shown to improve regression near the high energy edge of the region-of-interest (ROI) by extending the training data to include even larger energies beyond the ROI. Close to the small energy edge of the ROI, one may need to extrapolate the training data into the unphysical, negative energy region. The regressed values lying outside the ROI are discarded/ignored in the analysis and this results in lower efficiency. In contrast, our method does not result in regressed values being discarded.
Furthermore, we see that the MSE-based calibration corresponding to the log-uniform prior is generally more biased than the calibration corresponding to the uniform prior at intermediate energies.
@@ III.4 Resolution estimation
Furthermore, we found that for most values of $E_{\rm true}$, the MSE-based calibration obtains a full resolution that closely matches that from the NF. However, it is unable to provide per-shower resolutions. The large full resolution of the MSE-based calibration at $E_{\rm true}=90$ GeV is due to edge effects of the $E_{\rm inc}$ training data distribution.
@@
Since the NF is able to access the complete likelihood, this enables it to capture non-Gaussian [^8] In principle, the Gaussian Ansatz [17, 18] should also be able to capture non-Gaussian components of the resolution. However, this has yet to be shown in a worked example. Previous works [18, 111] only attempted to compute the Gaussian resolution. Also, the Gaussian Ansatz is trained in a different way (see Ref. [17]) compared to the NF and is more sensitive to hyperparameter choices. components of the resolution. We quantify this by defining an asymmetry observable based on the upper and lower bounds of the 68% confidence interval of the MLE prediction. In particular, the asymmetry is computed as the ratio $\frac{\Delta_{L}}{\Delta_{U}}$, where $\Delta_{L}$ ($\Delta_{U}$) is the difference between the lower (upper) bound of the 68% confidence interval and the MLE.
@@ IV Conclusions
We demonstrate a second advantage of using MLE-based calibration methods over direct regression by estimating per-shower resolutions. The average estimated per-shower resolution obtained from CaloFlow closely aligns with the full resolution. In contrast, the direct regression approach yields only point estimates and lacks the capacity to make per-shower resolution predictions. For CaloFlow, we found that smaller per-shower resolutions tend to coincide with more accurately calibrated showers. This is evidence that the per-shower resolutions from CaloFlow are reliable and have the potential to be leveraged for improved calibration. Notably, utilizing a NF for calibration grants access to the complete resolution function. This characteristic enables the NF to capture asymmetries in the resolution. In future work, it would be interesting to study how the per-shower resolution information obtained by the NF can be utilized to further improve the calibration.
@@ Appendix A Architecture and training
Here we briefly describe the architecture and training procedure used for CaloFlow (see Refs. [64, 65] for more details). There are some differences compared to the implementation in the original CaloFlow papers [64, 65], but most of the main algorithm remains the same. One main difference is that only Flow-I is used in this study.
The flows used in this work are Masked Autoregressive Flows (MAFs) [113] with compositions of Rational Quadratic Splines (RQS) [114] as transformations. The RQS transformations are parametrized using neural networks known as MADE blocks [115]. Identical flow architectures are used in each of the two cases with uniformly and log-uniformly distributed $E_{\rm inc}$. Each flow consists of six MADE blocks, each with two hidden layers of 64 nodes. The RQS transformations are defined with 8 bins and a tail bound of 14.